OpenStudy (anonymous):

Shoppers at a local department store were asked to complete a survey of their shopping experience. The results are shown in the table below: Satisfied with Service Not Satisfied with Service Totals Made a Purchase: 130 494 624 Did Not Make a Purchase: 715 183 898 Totals: 845 677 1522 a. What is the probability that a shopper selected at random made a purchase? b. What are the odds that a shopper selected at random was satisfied with their purchase?

OpenStudy (anonymous):

please help!!

OpenStudy (anonymous):

Give me a sec.......

OpenStudy (anonymous):

It's been awhile but I think for A, you have to divide total number of people made a purchase by total people overall. For B, I think you have to divide just the people who are satisfied with their purchase and total number of people over all.

OpenStudy (anonymous):

ok.. ill try that. ill leave this open a few more minutes to see if anyone who knows for sure can chime in. otherwise i'll go with that. thanks..

OpenStudy (anonymous):

No problem, its been a while I've done probability, right now doing parametrics and other crazy stuff in calculus III.

OpenStudy (anonymous):

So, I got 0.41 for the first question and 0.56 for the second

OpenStudy (anonymous):

Ok, so 0.41 and 0.56 is in decimal form, to put it in percentage was, just times 100 to turn it into a percentage form.

OpenStudy (anonymous):

right. it doesn't specify it's supposed to be in percent form. you think it should be?

OpenStudy (anonymous):

Well usually probability means what is the chance of something happening, etc. So if you say it is a 0.53 chance of winning the lottery vs. a chance of 53% of winning the lottery. The percentage form just makes more sense.

OpenStudy (anonymous):

gotcha. thanks..

OpenStudy (kropot72):

The odds are the ratio of the probability of an event occurring to that of its not occurring. P(satisfied) = 845/1522 P(not satisfied) = 677/1522 Odds(satisfied) = (845/15220) : (677/1522) = 845 : 677

OpenStudy (anonymous):

i am confused. so it was wrong what we did before?

OpenStudy (kropot72):

Yes. The odds are a ratio of probabilities.

OpenStudy (anonymous):

so what's the actual answer?

OpenStudy (anonymous):

That's not what the question says For A, it says whats the probability that a shopper made a purchase\[\frac{ 624 }{ 1522 }=.4099*100=40.99percent\] For B, it says odds of satisfied WITH their purchase\[\frac{ 130 }{ 624 }:\frac{ 494 }{ 624 }=130:494\]

OpenStudy (kropot72):

The odds that a shopper selected at random was satisfied with their purchase is 845 : 677

OpenStudy (anonymous):

oh so b would be 26%

OpenStudy (anonymous):

But the 845 includes the ones that DID NOT make a purchase, it asks the odds of the ones who MADE a purchase.

OpenStudy (anonymous):

or should i put it as 130 : 494

OpenStudy (anonymous):

I think it would be 130:494 because it was the odds selected random for the ones who were satisfied with their purchase hence meaning that they MADE a purchase.

OpenStudy (anonymous):

i messed up it is service not purchase.. i just noticed when i went to put in the answer. so it should be 845 : 677

OpenStudy (anonymous):

alright.. im about to close this and give hwtrain best answer unless you have something to say kropt, i see you're typing

OpenStudy (anonymous):

but like i said, it doesnt say purchase i goofed.. it just says service. which means it should be 845 : 677

OpenStudy (anonymous):

Oh if thats case @kropot72 was right all along... I bring dishonorable to my family........

OpenStudy (anonymous):

haha i completely messed you guys up sorry

OpenStudy (anonymous):

alright looks like i am good. thanks for the help

OpenStudy (anonymous):

*sniffle sniffle*, I got a math problem wrong :(

OpenStudy (kropot72):

Just a point to note. The maximum value of probability is 1. In statistics the maximum value of probability would not usually be stated as 100%.